OpenStudy (loveandpeace):

the perimeter of a right triangle is 12 and 6 sq units.what would be the measure or all three sides

OpenStudy (anonymous):

can you tell which number is the hypotenuse and which is the side?

OpenStudy (loveandpeace):

no

OpenStudy (anonymous):

lemme see what i can do.

OpenStudy (loveandpeace):

thanks

OpenStudy (akashdeepdeb):

You have got 3 equations! Let a and b be the right hand sides! and c be the hypotnuse! So a+b+c = 12 a*b/2 = 6 a2 + b2 = c2 [Pythagorus theorem] Solve the simultaneous equation now! :)

OpenStudy (loveandpeace):

ok thanks

OpenStudy (akashdeepdeb):

:)